OpenStudy (anonymous):

Find the probability of rolling an even number and then rolling a 6 when a number cube is rolled twice

OpenStudy (anonymous):

3

OpenStudy (anonymous):

(2,6) (6,2) (4,6) (6,4) (6,6) so there's 5 of them out of 36

OpenStudy (anonymous):

oh wait, the order matters so, there's only 3 of them. P = 3/36

OpenStudy (anonymous):

But there is only one die

OpenStudy (anonymous):

It's rolled TWICE

OpenStudy (anonymous):

Oh ok thank you

OpenStudy (anonymous):

It wasn't right

OpenStudy (anonymous):

3/36 = 1/12